OpenStudy (zeesbrat3):

A sample of gas occupies a volume of 74.7 mL. As it expands, it does 123.3 J of work on its surroundings at a constant pressure of 783 torr. What is the final volume of the gas?

OpenStudy (owen3):

Are you familiar with the gas laws?

OpenStudy (zeesbrat3):

I am not too strong with them

OpenStudy (owen3):

work= -PΔV -123.3 J = -783 torr × (Vfinal - 74.7 mL)

OpenStudy (zeesbrat3):

So, \[.157 = (Vfinal - 74.6 mL)\]

OpenStudy (zeesbrat3):

\[Vfinal = 74.757\] ?

OpenStudy (owen3):

We have to make some conversions. Torr is not an SI unit.

OpenStudy (zeesbrat3):

What would we convert it to?

OpenStudy (zeesbrat3):

Right!

OpenStudy (owen3):

we need the units on the left to match the units on the right. so we can convert Joules to L atm.

OpenStudy (owen3):

work= -PΔV -123.3 J = -783 torr × (Vfinal - 74.7 mL) -123.3 × (1L atm)/(101.32 J) = -783 torr × (1atm/760 torr) × (Vfinal -74.7 ml × 1L/1000 mL)

OpenStudy (zeesbrat3):

\[1.21694 = -1.03(Vfinal - .0747)\]

OpenStudy (zeesbrat3):

Sorry, the first one is negative

OpenStudy (zeesbrat3):

\[-1.21694=−1.03(Vfinal−.0747)\]

OpenStudy (zeesbrat3):

\[10.89 = (Vfinal - .0747)\] \[Vfinal = 10.965\]

OpenStudy (zeesbrat3):

But then we have to convert back to mL, so would it be 10965 mL?

OpenStudy (owen3):

doing this on paper, one minute

OpenStudy (owen3):

You should not round intermediate calculations. Ideally put it in one equation. -123.3 /101.3 = -783 / 760 ( v - 74.7/1000)

OpenStudy (zeesbrat3):

I did too when I did it the second time. I put 12.1694 by accident into my calculator

OpenStudy (zeesbrat3):

Gotcha, that way it is more accurate

OpenStudy (owen3):

I got 1.255889902 L or 1255.889902 mL which is about 1255.9 mL

OpenStudy (zeesbrat3):

Thank you for your help. I'm looking at another problem and it actually makes sense now!

OpenStudy (owen3):

Great :)

OpenStudy (zeesbrat3):

Again, thank you for teaching

OpenStudy (owen3):

your welcome